A rectangular solid has a base with length 5 cm and width 2 cm. If the volume of the solid is 100 cm3, find the height of the solid. [Hint: The volume of a rectangular solid is given by V = LWH.]
Need help trying to figure this out
Thank you

Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

Just substitute the respective variables into your formula v = lwh and solve away.
100 = 5*2*h, or 100 = 10h, or h = 100/10 = 10 cm., the height of the solid.

Answer by mananth(16946)

Vlome of the solid = 100 cm^3
..
Length = 5 cm
Width = 2cm
height = x
..
Volume V = L*W*H
V=5*2*x
5*2*x=100
10*x=100
/10
x= 10 cm the height
...
m.ananth@hotmail.ca